Status Report

ISS On-Orbit Status 5 Mar 2002

By SpaceRef Editor
March 5, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously or below.

After wake-up at the regular 1:00 am EST, the ISS crew began another busy day of Week 12, focusing on a variety of systems tasks and particularly on an 8A procedures review, using uplinked background material [while “over” at the Hubble telescope Jim Newman and Mike Massimino finished the second EVA (7h 16min) with successful installation of the second solar array, its associated electronics components and the RWA (reaction wheel assembly), all working flawlessly on ground testing].

Throughout the day, crewmembers performed noise measurements in the ISS with acoustic dosimeters, a periodic survey. SLM (sound level meter) readings were then transferred to the computer for later downlinking.

FE-1 Carl Walz set up the EARTHKAM at the Lab window and its power and computer connections for another run, remotely operated by participating school students on the ground taking selected Earth imagery. After the installation, EARTHKAM was activated.

As part of regular Vozdukh maintenance, CDR Yuri Onufrienko performed a scheduled test of the Vozdukh emergency vacuum valve (AVK). The CO2 removal system was then recycled and is now again working in Manual mode 5.

Onufrienko and Walz worked on the required relocation of cargo bundles temporarily stowed in the FGB and held a tagup via S-band with ground specialists to discuss the activity.

The two flight engineers completed tone tests of the C&W (caution and warning) system.

After yesterday’s ARIS (active rack isolation system) R&R, Dan Bursch today worked on ARIS pushrod calibration, to adjust the upper actuators for proper pushrod positioning. This got payload operators at POIC (Payload Operations Integration Center) one step closer to getting ARIS up and running again. EXPRESS Rack 2 had been powered up, and ARIS was up and in idle mode for the calibration activity. Additional calibration work is planned for tomorrow.

The planned reconfiguration of ARIS ICE (ARIS isolation characterization experiment) by Carl Walz for POP (payload operations processor) control by disconnecting a shorting plug and connecting a wiring harness was deferred until Bursch’s calibration activity is complete.

CDR Onufrienko completed the daily routine tasks of SOSH life support inspection/maintenance and preparing the IMS (inventory management system) “delta” file. He also tagged up with a ground specialist on the counter readings of the SVO water supply system collected yesterday.

All crewmembers performed their daily physical exercise program.

After today’s review by the crew of the uplinked “big picture” instructions for the upcoming robotics work, an all-discipline 8A overview conference is planned for tomorrow, followed by a robotics specific DOUG (dynamic operations ubiquitous graphics) review of the SSRMS operations planned for 3/7 (Thursday) and a crew conference. On Thursday, an S0 dry run will be conducted, supported by the ground by 8A astronaut Ellen Ochoa as MS2.

Tonight at 10:37pm and 11:28pm EST, Moscow will command a two-burn reboost of the ISS by Progress 6P thrusters via SM radiolink and the US-21 matching units (Russian tier 3 “slave” computers). Crew support is not required.

CEO (crew earth observation) target areas today were Typhoon Mitag (Philippines), (Dynamic Event Site: Major Typhoon Mitag is bearing down on the northern Philippine Islands with 120-kt winds gusting to 140-kts. Of interest: documenting this storm’s cloud pattern and well-defined eye feature), New Guinea (NASA has no useful photos to date of the small ice fields and glaciers of the Snow Mountains of western New Guinea. Weather is nearly always poor. Crew was to look for any views of opportunity through cloud breaks using the ESC [electronic still camera] here), Ganges River Delta (most of this target region was to the right of track this pass. Crew to try for regional context shots of this large, populous delta system), E. Mediterranean Dust and Smog (satellite imagery suggests a large plume of Saharan dust spreading northeastward over the Aegean Sea and Turkey. Of interest: oblique shots either side of track to help assess the density and extent of this dust), European Smog (high pressure over the western Mediterranean is slowly breaking down this pass, but ISS crossed the NE coast of Spain, crew was to try oblique and limb shots to the left of track to detect aerosol buildup), Congo-Zimbabwe Biomass Burning (with only fair weather cloud patterns this pass, crew was to try for oblique views to detect fire plumes either side of track), Industrialized Southeastern Africa (high pressure dominates South Africa at this time. Of interest as ISS approached the Mozambique coast from the NW: oblique views to right of track along the coast with aerosol plumes moving out to sea there), Gulf of St. Lawrence (with clearing from the SW, crew was to look to the right of track this pass for ice accumulation in the vicinity of Prince Edward Island), Eastern United States (a we=”r storm blasted this area over the weekend. Crew to try for detailed context views of snowfall patterns and aerosol output either side of track), Canadian Rocky Mountains (with most of this target to the left of track this pass, crew to try for detailed context views of snow cover and terrain features in oblique views), and Parana River (this pass could provide detailed context shots of much of the Parana River valley as harvest season begins there. Of interest: oblique views either side of track).

U.S. and Russian Segment Status (as of 2:13 pm EST today):

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On (16 Amps mode, the lowest possible setting). Vozdukh CO2 scrubber is ON in MANUAL cycle mode #5 (vacuum pump failed). U.S. CDRA CO2 scrubber is in Standby.
  • BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, #2 in Purify mode.
  • SM Working Compartment: Pressure (mmHg) — 758, temperature (deg C) — 26.5, ppO2 (mmHg) — 149.0, ppCO2 (mmHg) — 3.8.
  • SM Transfer Compartment: Pressure (mmHg) — 762, temperature (deg C) — 21.5; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • FGB Cabin: Pressure (mmHg) — 756, temperature (deg C) — 22.0; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Node: Pressure (mmHg) — 749.61, temperature (deg C) — 23.8 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab: Pressure (mmHg) — 751.58, temperature (deg C) — n/a, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
  • Joint Airlock (Equip. Lock): Pressure (mmHg) — 751.28, temperature (deg C) — n/a; shell heater temp (deg C) — 21.9, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1: Shell heater temp (deg C) — 23.7.
  • PMA-2: Shell heater temp (deg C) — 12.2.

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment (USOS) not available because MCA [major constituent analyzer] is failed and in Extended Life mode [= a state that preserves mass spectrometer vacuum but produces no pp data]. Russian GA (gas analyzer) readings of ppCO2 in the SM are invalid).

Electrical Power Systems (EPS):

  • Beta Gimbal Assembly (BGA) 2B in Autotrack mode, BGA 4B in Autotrack mode (during XPOP).
  • SM batteries: battery #7 is cycling; all other batteries (7) in “partial charge” mode.
  • FGB battery #5 is offline; battery #1 is cycling; all other batteries (4) are in “partial charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Thermal Control Systems:

  • Air conditioner SKV-1 is Off (Freon leak). SKV-2 is On.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby, all with new R2 s/w.
  • GNC-2 MDM is prime; GNC-1 is back-up, with R1 s/w.
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational.

Communications Systems:

  • S-band is operating nominally.
  • Ku-band is operating nominally in open loop pointing mode.
  • Audio subsystem operating nominally.
  • Video subsystem operating nominally.
  • MCOR (medium-rate communications outage recorder) operating nominally.


  • SSRMS/Canadarm2 at Port Stow position, with Keep Alive power on both strings.
  • RWS (robotics workstations) are Off.

ISS Orbit (as of this afternoon, 3:10pm EST):

  • Mean altitude — 381.5 km
  • Apogee — 388.2 km
  • Perigee — 374.7 km
  • Period — 92.2 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0010032
  • Orbits per 24-hr. day — 15.62
  • Altitude decrease — 350 m (mean) in last 24 hours
  • Solar Beta Angle: -39.7 deg (magnitude decreasing)
  • Revolutions since FGB/Zarya launch (Nov. ’98) — 18791
  • Current Flight Attitude — XPOP (x-axis perpendicular to orbit plane [yaw: ~0 deg, pitch: 5 deg., roll: 0 deg])

For more on ISS orbit and naked-eye visibility dates/times, see

SpaceRef staff editor.